OpenStudy (anonymous):

solve the following system of equations x+9y=6 (1) x=5-9y (2)

OpenStudy (anonymous):

x+9y=6 & x=5-9y (5-9y)+9y=6 5=6 So there are no interceptions.

OpenStudy (anonymous):

but what about the 1 and 2 you did not mention them?